OpenStudy (josedavid):

please help me Simplify. square root of 5 over cube root of 5 5 to the power of negative 1 over 6 5 to the power of negative 1 over 6 5 to the power of 5 over 6

OpenStudy (jusaquikie):

did you write all the options right? I think you should have a (5)^(1/6) option not two (5)^(-1/6) options

OpenStudy (jusaquikie):

\[\frac{ \sqrt{5} }{ \sqrt[3]{5} } = \frac{ \left( 5^{1/2} \right) }{ \left( 5^{1/3} \right) }=\frac{ \left( 5^{3/6} \right) }{\left( 5^{2/6} \right) }= 5^{1/6}=\sqrt[6]{5}\]